Fermi National Laboratory
 

Operation Summary for October 21, 2011 - October 24, 2011

 

NuMI reported receiving 8.27E18 protons on target for the period from 10/17/11 to 10/24/11.

Friday October 21

The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, with the Neutron Therapy Facility (NTF) preparing to treat a patient later in the morning, and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, a MI person began conducting some aperture scans of the 602 area at 9:51 AM. Operations established beam to the Meson FTBF experiment T-1008. Operators tuned the MI phase injection system at 3:14 PM. Booster and Antiproton Source (Pbar) personnel conducted studies.

During the evening shift, Booster personnel completed their study around 5:46 PM. Pbar Accumulator studies ended 7:43 PM. Operations disabled beam to T-1008 at midnight.

Saturday October 22

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, a MI RF station (MIRF14) tripped off twice in about four hours; it reset.

During the day shift, Operations established beam to T-1008 at 10:26 AM. At 11:38 AM, MIRF17 tripped off; operators investigated, reported that there was something wrong with the bias power supply, bypassed the station, and contacted a RF person. At 1:17 PM, BRF2 tripped off on an overcurrent indication; operators lowered the Booster intensity for about 40 minutes until a RF person came in and reset the bias power supply.

During the evening shift, at 8:32 PM, the P1 transfer line suffered a vacuum burst; it recovered.

Sunday October 23

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and T-1008.

During the midnight shift, operators disabled beam to T-1008 at 12:17 AM. At 3:51 AM, LRF5 tripped off for the first of five times in about 20 minutes; it reset.

During the day shift, Operations established beam to T-1008 at 10:06 AM.

During the evening shift, at 5:43 PM, operators reported that the H- Source, the backup source, was having an arc voltage problem; a Linac person came in and stabilized the source. Operations disabled beam to T-1008 at 10:02 PM.

Monday October 24

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations monitored the Accelerator Complex.

Operations Coordinator (OPSCO)

The OpsCo approved of the plans by MI and Pbar personnel to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/24/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory